OpenStudy (hitaro9):

Is there a way to indicate the # of primes in the prime factorization of a number?

OpenStudy (hitaro9):

ie n=(p1 ^a1)(p2^a2)...(pk^ak) is there a function that returns k?

OpenStudy (adi3):

wht grade is tht

OpenStudy (hitaro9):

Number theory

OpenStudy (hitaro9):

College

OpenStudy (adi3):

wooow, i m just in 9th grade.. lol

OpenStudy (adi3):

any way i'll try to answer ur question

OpenStudy (hitaro9):

Haha it's fine. I don't think it's anything too complicated for someone's who taken the course, or even someone who's particularly good at google. I'm just drawing a blank

OpenStudy (adi3):

http://en.wikipedia.org/wiki/Fundamental_theorem_of_arithmetic is this the stuff

OpenStudy (adi3):

can u check

OpenStudy (hitaro9):

Right, I need a function that returns the # of primes in one of those lists though

OpenStudy (adi3):

is it similar to ur stuffff

OpenStudy (hitaro9):

Like, say 9 has prime factorization 3x3, I would need f(9) = 2 or say 27 has prime factorisation 3x3x3, I would need f(27) =3

OpenStudy (hitaro9):

It's in the general vicinity yeah

OpenStudy (adi3):

i know wht does f(x) mean

OpenStudy (hitaro9):

I need f(x) = number of prime factorizations

OpenStudy (hitaro9):

Whatever that function would be

OpenStudy (hitaro9):

I need that as a step in my homework, but I can't figure out how to get that notation. I suspect it would be some manipulation of the Euler-Phi function

OpenStudy (adi3):

@iGreen

OpenStudy (adi3):

@ribhu

OpenStudy (adi3):

i'm tryna invite people so they can help u

OpenStudy (hitaro9):

Yeah I see thanks. I guess if they knew they'd pop in and help anyways haha

OpenStudy (adi3):

wait a while

OpenStudy (adi3):

some one came

OpenStudy (hitaro9):

@iGreen I'm studying Mobius inversion, Euler phi functions, Mersenne primes etc. I can't figure out some manipulation to give a function that returns exactly the number of prime divisors of a function.

OpenStudy (hitaro9):

Okay I gotta commute to class. If anyone sees this while I'm gone please feel free to answer, I'll look at it before class and give medal c:

OpenStudy (adi3):

dude can u just send me a message for a sec

OpenStudy (adi3):

@Hitaro9

OpenStudy (rational):

what you're looking for doesn't exist, but i think what you need is the function \(\omega\) see http://mathworld.wolfram.com/DistinctPrimeFactors.html

OpenStudy (kainui):

Just a quick thing to add on cause I kept forgetting but this helped me remember these functions better. Capital Omega represents the total number of prime factors and lowercase omega returns the number of distinct primes. So we can set up this nice inequality that is cute cause it's suggestive based on just the size of the letters haha. \[\Large \Omega(n) \ge \omega(n)\] So for a quick example: \[\Large \omega(45)=\omega(3^25^1)=2 \\ \Large \Omega(45)=\Omega(3^25^1)=3\] Too bad there's not really any way to evaluate these except by factoring a number. :/